2x^2 – 7x + 3 = 0
roots finddddddfjhbvuejnbvtbverjbnvrjbvrhbv rhbvr

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

On comparing the given equation
ax^2 + bx + c = 0,
a = 2, b = -7 and c = 3
By using quadratic formula,
⇒ x = (7±√(49 – 24))/4
⇒ x = (7±√25)/4
⇒ x = (7±5)/4
⇒ x = (7+5)/4 or x = (7-5)/4
⇒ x = 12/4 or 2/4
∴ x = 3 or ½

Thanks